### v2

# Anderson-Picard based nonlinear preconditioning of the Newton iteration for non-isothermal flow simulations

Abstract
We propose, analyze, and test a nonlinear preconditioning technique to improve the Newton iteration for non-isothermal flow simulations. We prove that by first applying an Anderson accelerated Picard step, Newton becomes unconditionally stable (under a uniqueness condition on the data) and its quadratic convergence is retained but has less restrictive sufficient conditions on the Rayleigh number and initial condition’s accuracy. Since the Anderson-Picard step decouples the equations in the system, this nonlinear preconditioning adds relatively little extra cost to the Newton iteration (which does not decouple the equations). Our numerical tests illustrate this quadratic convergence and stability on multiple benchmark problems. Furthermore, the tests show convergence for significantly higher Rayleigh number than both Picard and Newton, which illustrates the larger convergence basin of Anderson-Picard based nonlinear preconditioned Newton that the theory predicts.
1 Introduction
The non-isothermal flow simulations are a multiphysics model of natural convection, and are commonly used to model flows with non-constant density or temperature. We consider this system, often called the Boussinesq system, on a finite connected domain $\Omega\subset\mathbb{R}^{d}$ $(d=2,3)$ with $\partial\Omega=\Gamma_{1}\cup\Gamma_{2}$, and is given by
$\begin{cases}u\cdot\nabla u+\nabla p-\nu\Delta u&=f+R_{i}(0~{}T)^{T}\text{ in }\Omega,\\ \nabla\cdot u&=0\text{ in }\Omega,\\ u\cdot\nabla T-\kappa\Delta T&=g\text{ in }\Omega,\\ \end{cases}$ (1)
with boundary conditions
$\begin{cases}u=0&\text{ on }\partial\Omega,\\ T=0&\text{ on }\Gamma_{1},\\ \nabla T\cdot n=0&\text{ on }\Gamma_{2}.\end{cases}$ (2)
Here, $u$ is the fluid velocity, $p$ is the pressure, $T$ is the temperature (or density), $\nu>0$ is the kinematic viscosity of the fluid, $\kappa>0$ is the thermal diffusivity, $f$ is the external forcing term, and $R_{i}>0$ is the Richardson number which accounts for the gravitational force and thermal expansion of the fluid. This model can be observed in nature such as in atmospheric models and katabatic winds, and is also observed in industry for natural ventilation, dense gas dispersion, insulation with double pane window, and solar collectors [6].
The boundary conditions we state above are no-slip for velocity and mixed homogenous Dirichlet and homogenous Neumann for temperature (the latter of which corresponds to perfect insulation), but our results are extendable to most common boundary conditions. Important physical constants are the Reynolds number $Re$, Prandtl number $Pr$ which describes the ratio of kinematic viscosity and thermal diffusivity, and the Rayleigh number $Ra$:
$Re=\nu^{-1}~{}~{}~{}~{}Pr=\frac{\nu}{\kappa}~{}~{}~{}~{}Ra=Ri~{}Re^{2}~{}Pr.\$
The non-isothermal flow system (1)-(2) is known to be well-posed under a small data condition [3], however uniqueness is lost for high $Ra$.
The Picard iteration and Newton iteration are common iterative methods for solving the system (1)-(2). The Picard iteration for this system is given by
$\begin{cases}u^{k}\cdot\nabla u^{k+1}+\nabla p^{k+1}-\nu\Delta u^{k+1}&=f+R_{i}(0~{}T^{k+1})^{T},\\ \nabla\cdot u^{k+1}&=0,\\ u^{k}\cdot\nabla T^{k+1}-\kappa\Delta T^{k+1}&=g,\end{cases}$ (3)
with boundary conditions
$\begin{cases}u^{k+1}=0&\text{ on }\partial\Omega,\\ T^{k+1}=0&\text{ on }\Gamma_{1},\\ \nabla T^{k+1}\cdot n=0&\text{ on }\Gamma_{2},\end{cases}$ (4)
and is known to admit stable solutions for any problem data (see Section 2). Furthermore, provided the data satisfies a smallness condition, the solution to Picard is unique and the iteration convergences linearly for any initial guess (see Section 2). Note that the system (3) decouples the equations because the heat transport equation is independent of velocity. While this all makes Picard very attractive with respect to efficiency at each iteration, the linear convergence is a potential drawback [15, 10, 20].
The Newton iteration for the system (1)-(2) is given by
$\begin{cases}u^{k+1}\cdot\nabla u^{k}+u^{k}\cdot\nabla u^{k+1}+\nabla p^{k+1}-\nu\Delta u^{k+1}&=f+R_{i}(0~{}T^{k+1})^{T}+u^{k}\cdot\nabla u^{k},\\ \nabla\cdot u^{k+1}&=0,\\ u^{k+1}\cdot\nabla T^{k}+u^{k}\cdot\nabla T^{k+1}-\kappa\Delta T^{k+1}&=g+u^{k}\cdot\nabla T^{k},\end{cases}$ (5)
with boundary conditions (4). This iteration will be shown in Section 2 to be well-posed and converge quadratically, but only for a very good initial guess and sufficiently small data (the sufficient conditions are more restrictive than for Picard). Newton is also more computationally expensive at each iteration compared to Picard since it does not decouple, and so one must solve fully coupled linear systems at each iteration.
The purpose of this paper is to consider the iteration defined by, at each step, first applying Anderson accelerated Picard and then applying Newton (i.e. Anderson-Picard based nonlinearly preconditioned Newton). Our intent of combining the iterations in this way is to unite the speed of Newton with the stability and robustness of Picard. This method is simple to implement and we show it has an improved robustness compared to Newton (both with respect to stability and the size of the convergence basin), but still converges quadratically. This work thus fits a theme of recent and fundamentally important research on nonlinear preconditioning [5, 14, 13].
The general idea of applying Anderson-Picard as a preconditioner was first done in 2024 when it was applied to the Navier-Stokes Equations (NSE) [17]. The purpose of this paper is to extend this important yet easy to implement idea to non-isothermal flow. The difference between NSE and non-isothermal flow simulations is foremost that it is coupled with another equation which incorporates temperature (or density) into the model. This is further complicated by the extra nonlinear terms that arise in this coupled equation. This consequently gives rise to extra difficulties in the analysis and computation, especially for iterative methods like Newton where the equations remain fully coupled. In other words, it is not at all a straightforward extension. Moreover, the analysis provided in this paper will reveal that the convergence basin for Anderson-Picard-Newton is dependent upon only the initial conditions for velocity. Note that for non-preconditioned Newton it depends on both the initial conditions of velocity and temperature (or density). This is concurrent with the accuracy of Anderson-Picard-Newton depending more critically on velocity in a similar manner.
We show herein analytically and numerically that the (Anderson-)Picard based nonlinearly preconditioned Newton, shortened to (AA-)Picard-Newton, converges quadratically, with better convergence properties than Newton as well as improved stability. We restrict our analysis to Picard-Newton, i.e no Anderson, and note that the results are extendable to AA-Picard-Newton with some extra technical difficulties. We also provide numerical results for both AA-Picard-Newton and Picard-Newton that show a significant improvement compared to Newton alone.
This paper is arranged as follows. We provide analysis of the well-posedness, stability, and convergence of Picard and Newton iterations in section 2. While much of this is known, the proofs and results will allow for a better comparison and analysis of Picard-Newton. In section 3 we give analysis on the stability and convergence of Picard-Newton. Following this in section 4, we provide numerical results for both AA-Picard-Newton and Picard-Newton applied to some benchmark tests.
2 Preliminaries
[…]
$X$ =$=\{v\in H^{1}(\Omega):v=0\text{ on }\partial\Omega\},$
$Q$ =$=\{q\in L^{2}(\Omega):\int_{\Omega}q=0\},$
$D$ =$=\{w\in H^{1}(\Omega):w=0\text{ on }\Gamma_{1}\text{ and }\nabla w\cdot n=0\text{ on }\Gamma_{2}\},$
$V$ =$=\{v\in X:(\nabla\cdot v,q)=0~{}\forall q\in Q\}.$
Let $X^{\ast}$, $V^{\ast}$, and $D^{\ast}$ be the dual spaces of $X$, $V$, and $D$, respectively. With slight abuse of notation, we also use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote the dual pairing of $X$ and $X^{\ast}$, $V$ and $V^{\ast}$, and $D$ and $D^{\ast}$. Then for example, the dual norm of $X$ is defined by
$\|F\|_{X^{\ast}}:=\sup_{z\in X}\frac{(F,z)}{\|z\|_{X}}.$
We also recall the Poincaré inequality for $z\in X\text{ or }D$, given by
$\|z\|\leq C_{p}\|\nabla z\|,$
where $C_{p}>0$ is a constant depending only on the size of the domain $\Omega$.
Define the trilinear functionals $b:X\times X\times X\rightarrow\mathbb{R}$ and $\hat{b}:X\times D\times D\rightarrow\mathbb{R}$ by :
$b(u,v,w)$ $:=(u\cdot\nabla v,w)+\frac{1}{2}((\nabla\cdot u)v,w),$
$\hat{b}(u,v,w)$ $:=(u\cdot\nabla s,t)+\frac{1}{2}((\nabla\cdot u)s,t).$
Then $b(u,v,w)$ and $\hat{b}(u,s,t)$ are skew-symmetric, i.e.,
$b(u,v,v)=0\text{ and }\hat{b}(u,s,s)=0.$
Furthermore, if $u$ is divergence free, then $b(u,v,w)=(u\cdot\nabla v,w)$, and thus $(u\cdot\nabla v,v)=0$. A well known bound for $b$ and $\hat{b}$ that is used herein is [12]: $\forall u,v,w\in X$, $s,t\in D$, $\exists C_{s}>0$ depending only on $|\Omega|$ such that
$|b(u,v,w)|$ $\leq C_{s}\|\nabla u\|\|\nabla v\|\|\nabla w\|,$ (6)
$|\hat{b}(u,s,t)|$ $\leq C_{s}\|\nabla u\|\|\nabla s\|\|\nabla t\|.$ (7)
The use of $b$ and $\hat{b}$ will allow for $X,Q,D$ to be replaced by finite dimensional subspaces and have the exact same analysis of the paper hold.
2.1 Anderson-Acceleration
Anderson Acceleration (AA) is an extrapolation technique which utilizes previously computed iterates to construct the next solution. It is known to improve convergence for linearly convergent fixed point iterations like Picard [18, 8, 16, 19, 11, 8]. Therefore, it is a natural to precondition Newton with Anderson-Picard.
AA is defined as follows [18]. Let $g:Y\rightarrow Y$ be a fixed point operator for a Hilbert space $Y$.
Algorithm 2.1.
AA for depth $m\geq 0$
Step 0: Choose $x_{0}\in Y$
Step 1: Find $w_{1}\in Y$ such that $w_{1}=g(x_{0})-x_{0}$. Set $x_{1}=x_{0}+w_{1}$.
Step k: For $k=2,3,...$. Set $m_{k}=\min\{k-1,m\}$.
  [a.] Find $w_{k}=g(x_{k-1})-x_{k-1}$.
  [b.] Solve the minimization problem for $\{\alpha_{j}^{k}\}_{k-m_{k}}^{k-1}$
$\min\|\left(1-\sum_{j=k-m_{k}}^{k-1}\alpha_{j}^{k}\right)w_{k}+\sum_{j=k-m_{k}}^{k-1}\alpha_{j}^{k}w_{j}\|_{Y}.$
  [c.] Set
$x_{k}=(1-\sum_{j=k-m_{k}}^{k-1}\alpha_{j}^{k})(x_{k-1}+w_{k})+\sum_{j=k-m_{k}}^{k-1}\alpha_{j}^{k}\alpha_{j}^{k}(x_{j-1}+w_{j})$
where $w_{j}=g(x_{j-1})-x_{j-1}$ may be referred to as the update step or also as the nonlinear residual. Setting $m=0$ returns the original fixed point iteration.
2.2 Non-isothermal flow background
The weak form of the non-isothermal flow system (1)-(2) is: Given $f\in X^{\ast}$ and $g\in D^{\ast}$, find ($\text{find }(u,p,T)\in(X,Q,D)\text{ satisfying}$ $\forall(v,q,w)\in(X,Q,D)$,
$\begin{cases}b(u,u,v)+(\nabla p,v)+\nu(\nabla u,\nabla v)&=(f,v)+R_{i}((0~{}T)^{T},v),\\ (\nabla\cdot u,q)&=0,\\ \hat{b}(u,T,w)+\kappa(\nabla T,\nabla w)&=(g,w).\end{cases}$
The spaces $(X,Q)$ satisfy an inf-sup condition, and thus the weak formulation is equivalently simplified using the solution space $V$: Find $(u,T)\in(V,D)$ satisfying $\forall(v,w)\in(V,D)$,
$\begin{cases}b(u,u,v)+\nu(\nabla u,\nabla v)&=(f,v)+R_{i}((0~{}T)^{T},v),\\ \hat{b}(u,T,w)+\kappa(\nabla T,\nabla w)&=(g,w).\end{cases}$ (8)
Lemma 2.1.
Any solution to the non-isothermal flow system (8) satisfies the a priori estimate
$\|\nabla T\|$ $\leq\kappa^{-1}\|g\|_{D^{\ast}}=:M_{2},$ (9)
$\|\nabla u\|$ $\leq\nu^{-1}\|f\|_{V^{\ast}}+R_{i}C_{p}^{2}\nu^{-1}M_{2}=:M_{1}.$ (10)
Proof.
We let $v=u$ and $w=T$ in (8) then using skew-symmetry gives us
$\begin{cases}\nu\|\nabla u\|^{2}=Ri((0~{}T)^{T},u)+(f,u),\\ \kappa\|\nabla T\|^{2}=(g,T).\end{cases}$ (11)
We upper bound the right hand side terms using the dual space norms, Cauchy-Schwarz, and Poincaré, which yields
$Ri((0~{}T)^{T},u)$ $\leq C_{p}^{2}Ri\|\nabla T\|\|\nabla u\|,$
$(f,u)$ $\leq\|f\|_{V^{\ast}}\|\nabla u\|,$
$(g,T)$ $\leq\|g\|_{D^{\ast}}\|\nabla T\|.$
Using these bounds in (11) and reducing provides us with
$\begin{cases}\|\nabla u\|\leq\nu^{-1}\|f\|_{V^{\ast}}+R_{i}C_{p}^{2}\nu^{-1}\|\nabla T\|,\\ \|\nabla T\|\leq\kappa^{-1}\|g\|_{D^{\ast}},\end{cases}$
and using the second of these bounds in the first,
$\|\nabla u\|\leq\nu^{-1}\|f\|_{V^{\ast}}+R_{i}C_{p}^{2}\nu^{-1}\kappa^{-1}\|g\|_{D^{\ast}}.$
This proves the result. ∎
The proof gives the bounds (9) and (10) for $T$ and $u$ respectively, and will be used throughout the paper for solutions to the non-isothermal flow system. These results are similar to those found in [3], and depending on how the problem is formulated (e.g. Richardson number $\neq 1$ or Prandtl number $\neq 1$), the constants can differ but the results are equivalent. It is known that (8) admits solutions for any $Ra>0$ [3], and this can be proven in the same way as the steady Navier-Stokes existence proof using the Leray-Schauder theorem [12]. Uniqueness of (8), however, requires a smallness assumption on the data. Below is a sufficient condition on the problem data that produces uniqueness.
Lemma 2.2.
Let $\alpha_{1}=C_{s}\nu^{-1}M_{1}$ and $\alpha_{2}=C_{s}\kappa^{-1}M_{2}$. If $C_{p}^{2}\nu^{-1}R_{i},\alpha_{2}+\alpha_{1}<1$, then solutions to (8) are unique.
Proof.
Supposing two solutions $(u_{1},T_{1})$ and $(u_{2},T_{2})$ to (8) exist, define $e_{u}=u_{1}-u_{2}$ and $e_{T}=T_{1}-T_{2}$. Now subtracting the systems with these two solutions gives $\forall v\in V,w\in D$,
$\begin{cases}b(u_{1},e_{u},v)+b(e_{u},u_{2},v)+\nu(\nabla e_{u},\nabla v)&=Ri((0~{}e_{T})^{T},v),\\ \hat{b}(u_{1},e_{T},w)+\hat{b}(e_{u},T_{2},w)+\kappa(\nabla e_{T},\nabla w)&=0.\end{cases}$
Taking $v=e_{u}$ and $w=e_{T}$ vanishes two nonlinear terms and leaves
$\begin{cases}\nu\|\nabla e_{u}\|^{2}&=Ri((0~{}e_{T})^{T},e_{u})-b(e_{u},u_{2},e_{u})\\ &\leq C_{p}^{2}R_{i}\|\nabla e_{T}\|\|\nabla e_{u}\|+C_{s}\|\nabla e_{u}\|^{2}\|\nabla u_{2}\|,\\ &\\ \kappa\|\nabla e_{T}\|^{2}&=-\hat{b}(e_{u},T_{2},e_{T})\\ &\leq C_{s}\|\nabla T_{2}\|\|\nabla e_{u}\|\|\nabla e_{T}\|.\end{cases}$
Next, using the bounds (10) and (9) and simplifying gives
$\begin{cases}\|\nabla e_{u}\|&\leq C_{p}^{2}\nu^{-1}R_{i}\|\nabla e_{T}\|+\alpha_{1}\|\nabla e_{u}\|,\\ \|\nabla e_{T}\|&\leq\alpha_{2}\|\nabla e_{u}\|.\end{cases}$ (12)
Adding these and simplifying results in
$(1-\alpha_{1}-\alpha_{2})\|\nabla e_{u}\|+(1-C_{p}^{2}\nu^{-1}R_{i})\|\nabla e_{T}\|\leq 0.\$
This provides the uniqueness of the velocity due to the assumption on the data. With this, uniqueness of the temperature follows immediately from the second bound in (12). ∎
2.3 Picard iteration
The weak formulations for Picard for the non-isothermal flow system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
$\begin{cases}b(u^{k},u^{k+1},v)+\nu(\nabla u^{k+1},\nabla v)&=(f,v)+R_{i}((0~{}T^{k+1})^{T},v),\\ \hat{b}(u^{k},T^{k+1},w)+\kappa(\nabla T^{k+1},\nabla w)&=(g,w).\end{cases}$ (13)
Note that the Picard iteration decouples the temperature equation and thus solving (13) is a two-step process where one first solves a scalar convection-diffusion problem and then an Oseen problem. Effective preconditioners for these linear systems exist in the literature [2, 7, 9, 4].
Lemma 2.3.
Any solution to the Picard iteration for the non-isothermal flow system satisfies the a priori estimate: for any $k=1,2...$,
$\|\nabla T^{k}\|\leq M_{2},$
$\|\nabla u^{k}\|\leq M_{1},$
Proof.
These results are proved analogously to those of Lemma 2.1. ∎
Lemma 2.4.
The Picard iteration (13) with data satisfying $\min\{1-\nu^{-1}\frac{C_{p}^{2}R_{i}}{2},1-\kappa^{-1}\frac{C_{p}^{2}R_{i}}{2}\}>0$, admits a unique solution.
Remark.
Due to Lemma 2.3 and since (13) is linear, solution uniqueness can be proven immediately. In the finite dimensional case, this will also imply existence of solutions.
Proof.
Let $Y=V\times D$ and at iteration $k+1$ define $A:Y\times Y\rightarrow\mathbb{R}$ and $F:Y\rightarrow\mathbb{R}$ by
$A((\hat{u},\hat{T}),(v,w)):$ $=b(u^{k},\hat{u},v)+\nu(\nabla\hat{u},\nabla v)+\hat{b}(u^{k},\hat{T},w)+\kappa(\nabla\hat{T},\nabla w)-R_{i}((0~{}\hat{T})^{T},v),$
$F((v,w))$ $=(f,v)+(g,w),$
so that the Picard iteration is given by $A((\hat{u},\hat{T}),(v,w))=F((v,w))$. Consider $A((\hat{u},\hat{T}),(v,w))$. Using (6), Cauchy-Schwarz, and Young’s inequality we lower bound the equation as
$A((\hat{u},\hat{T}),(\hat{u},\hat{T}))$ $=b(u^{k},\hat{u},\hat{u})+\nu\|\nabla\hat{u}\|^{2}+\hat{b}(u^{k},\hat{T},\hat{T})+\kappa\|\nabla\hat{T}\|^{2}-R_{i}((0~{}\hat{T})^{T},\hat{u})$
$\geq\nu\|\nabla\hat{u}\|^{2}+\kappa\|\nabla\hat{T}\|^{2}-\frac{C_{p}^{2}R_{i}}{2}\|\nabla\hat{T}\|^{2}-\frac{C_{p}^{2}R_{i}}{2}\|\nabla\hat{u}\|^{2}$
$\geq\min\{\nu-\frac{C_{p}^{2}R_{i}}{2},\kappa-\frac{C_{p}^{2}R_{i}}{2}\}\|(\hat{u},\hat{T})\|^{2}_{Y}.$
Hence $A$ is coercive. Continuity of $A$ and $F$ follow easily using the bounds and lemmas above. Thus Lax-Milgram applies and gives existence and uniqueness of (13).
We can now define the solution operator for the Picard iteration. Define $g_{P}:(V,D)\rightarrow(V,D)$ to be the solution operator for (13): $g_{P}(u^{k},T^{k})=(u^{k+1},T^{k+1})$. Because the solutions to Picard are unique, this is a well defined operator.
Lemma 2.5.
Consider the Picard iteration (13) with data satisfying $C_{p}^{2}\nu^{-1}R_{i}<1$ and $\alpha_{1}+\alpha_{2}<1$. Then the iteration converges linearly with rate $\alpha_{1}+\alpha_{2}$. In particular we have
$\|\nabla(T-T^{k+1})\|\leq\alpha_{2}\|\nabla(u-u^{k})\|,$
and
$\|\nabla(u-u^{k+1})\|\leq(\alpha_{1}+\alpha_{2})\|\nabla(u-u^{k})\|.$
Remark.
Note that the sufficient conditions for convergence of Picard are the same as the sufficient conditions for the uniqueness of solutions to the non-isothermal flow system.
Proof.
We subtract (8) from (13) and choose $v=e^{k+1}$ and $q=e_{T}^{k+1}$. Using skew-symmetry , vanishes two nonlinear terms and leaves the equality
[…]
Next we use Cauchy-Schwarz, Poincaré, and (6) to upper bound these equations as
$\begin{cases}\nu\|\nabla e^{k+1}\|^{2}&\leq C_{p}^{2}R_{i}\|\nabla e_{T}^{k+1}\|\|\nabla e^{k+1}\|+C_{s}\|\nabla e^{k}\|\|\nabla u\|\|\nabla e^{k+1}\|,\\ \kappa\|\nabla e_{T}^{k+1}\|^{2}&\leq C_{s}\|\nabla e^{k}\|\|\nabla T\|\|\nabla e_{T}^{k+1}\|.\end{cases}$
Then we reduce and apply Lemma 2.1 to get
$\begin{cases}\|\nabla e^{k+1}\|&\leq C_{p}^{2}\nu^{-1}R_{i}\|\nabla e_{T}^{k+1}\|+\alpha_{1}\|\nabla e^{k}\|,\\ \|\nabla e_{T}^{k+1}\|&\leq\alpha_{2}\|\nabla e^{k}\|.\end{cases}$
this gives the bound
$\|\nabla e_{T}^{k+1}\|\leq\alpha_{2}\|\nabla e^{k}\|,$
Adding the equations and reducing gives
$\|\nabla e^{k+1}\|+(1-C_{p}^{2}R_{i}\nu^{-1})\|\nabla e_{T}^{k+1}\|^{2}$ $\leq C_{s}(\nu^{-1}\|\nabla u\|+\kappa^{-1}\|\nabla T\|)\|\nabla e^{k}\|$
$\leq C_{s}(\nu^{-1}M_{1}+\kappa^{-1}M_{2})\|\nabla e^{k}\|$
$\leq(\alpha_{1}+\alpha_{2})\|\nabla e^{k}\|.$
Finally, using the assumptions on the data finishes the proof. ∎
2.4 Newton iteration
The weak formulation for the Newton iteration for the non-isothermal flow system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
$\begin{cases}b(u^{k+1},u^{k},v)+b(u^{k},u^{k+1},v)+\nu(\nabla u^{k+1},\nabla v)&=(f,v)+R_{i}((0~{}T^{k+1})^{T},v)\\ &~{}~{}~{}+b(u^{k},u^{k},v),\\ \hat{b}(u^{k+1},T^{k},w)+\hat{b}(u^{k},T^{k+1},w)+\kappa(\nabla T^{k+1},\nabla w)&=(g,w)+\hat{b}(u^{k},T^{k},w).\end{cases}$ (14)
Lemma 2.6.
Consider the Newton iteration (14) with data and solutions satisfying
$C_{s}\nu^{-1}\|\nabla(u^{k}-u)\|+\alpha_{1}+C_{s}\nu^{-1}\|\nabla(T^{k}-T)\|+C_{s}\nu^{-1}M_{2}+C_{p}^{2}R_{i}<1,$
$\frac{C_{s}\kappa^{-1}}{4}\|\nabla(T^{k}-T)\|+\frac{\alpha_{2}}{4}+\frac{C_{p}^{2}\kappa^{-1}}{4}R_{i})<1.$
Then there exists a unique solution to the iteration.
Proof.
We proceed just like for Picard and will use Lax-Milgram. Let $Y=V\times D$ and at iteration $k+1$ define $A:Y\times Y\rightarrow\mathbb{R}$ and $F:Y\rightarrow\mathbb{R}$ by
$A((\hat{u},\hat{T}),(v,w)):$ $=b(\hat{u},u^{k},v)+b(u^{k},\hat{u},v)+\nu(\nabla\hat{u},\nabla v)-R_{i}((0~{}\hat{T})^{T},v)+\hat{b}(\hat{u},T^{k},w)+\hat{b}(u^{k},\hat{T},w),$
$~{}~{}~{}+\kappa(\nabla T,\nabla w)$
$F((v,w))$ $=(f,v)+b(u^{k},u^{k},v)+(g,w)+\hat{b}(u^{k},T^{k},w).$
Then the Newton iteration is given by $A((\hat{u},\hat{T}),(v,w))=F((v,w))$. Continuity follows by applying (6), Poincaré, Cauchy-Schwarz, and Young’s inequality
$A((\hat{u},\hat{T}),(v,w))$ $=b(\hat{u},u^{k},v)+b(u^{k},\hat{u},v)+\nu(\nabla\hat{u},\nabla v)-R_{i}((0~{}\hat{T})^{T},v)$
$~{}~{}~{}+\hat{b}(\hat{u},T^{k},w)+\hat{b}(u^{k},\hat{T},w)+\kappa(\nabla\hat{T},\nabla w)$
$\leq(2C_{s}\|\nabla u^{k}\|+\nu)\|\nabla\hat{u}\|\|\nabla v\|+C_{p}^{2}R_{i}\|\nabla\hat{T}\|\|\nabla v\|$
$~{}~{}~{}+C_{s}\|\nabla\hat{u}\|\|\nabla T^{k}\|\|\nabla w\|+(C_{s}\|\nabla u^{k}\|+\kappa)\|\nabla\hat{T}\|\|\nabla w\|$
$\leq\sigma_{max}(\|\nabla\hat{u}\|+\|\nabla\hat{T}\|)(\|\nabla v\|+\|\nabla w\|)$
$\leq\sigma_{max}\|(\hat{u},\hat{T})\|_{Y}\|(v,w)\|_{Y},$
where $\sigma_{max}=\max\{(2C_{s}\|\nabla u^{k}\|+\nu),C_{p}^{2}R_{i},C_{s}\|\nabla T^{k}\|,C_{s}\|\nabla u^{k}\|+\kappa\}$. The continuity of $F$ is shown using the dual norms and (6) as
$F((v,w))$ $=(f,v)+b(u^{k},u^{k},v)+(g,w)+\hat{b}(u^{k},T^{k},w)$
$\leq\|f\|_{V^{\ast}}\|v\|+C_{s}\|\nabla u^{k}\|^{2}\|\nabla v\|+\|g\|_{D^{\ast}}\|\nabla w\|+C_{s}\|\nabla u^{k}\|\|\nabla T^{k}\|\|\nabla w\|$
$\leq(\|f\|_{V^{\ast}}+C_{s}\|\nabla u^{k}\|^{2}+\|g\|_{D^{\ast}}+C_{s}\|\nabla u^{k}\|\|\nabla T^{k}\|)\|(v,w)\|_{Y}.$
Next we show $A$ is coercive, and this will require a restriction on the problem data. Here we use Cauchy-Schwarz and skew symmetry to lower bound $A$ as
$A((\hat{u},T),(\hat{u},\hat{T}))$ $=b(\hat{u},u^{k},\hat{u})+b(u^{k},\hat{u},\hat{u})+\nu\|\nabla\hat{u}\|^{2}+\hat{b}(\hat{u},T^{k},T)+b(u^{k},\hat{T},\hat{T})+\kappa\|\nabla\hat{T}\|^{2}$
$~{}~{}~{}-R_{i}((0~{}\hat{T})^{T},\hat{u})$
$\geq b(\hat{u},u^{k},\hat{u})+\nu\|\nabla\hat{u}\|^{2}+\hat{b}(\hat{u},T^{k},\hat{T})+\kappa\|\nabla\hat{T}\|^{2}-C_{p}^{2}R_{i}\|\nabla\hat{T}\|\|\nabla\hat{u}\|$
$\geq b(\hat{u,}u^{k}-u,\hat{u})+b(\hat{u},u,\hat{u})+\nu\|\nabla\hat{u}\|^{2}+\hat{b}(\hat{u},T^{k}-T,\hat{T})$
$~{}~{}~{}+\hat{b}(\hat{u},T,\hat{T})+\kappa\|\nabla\hat{T}\|^{2}-C_{p}^{2}R_{i}\|\nabla\hat{T}\|\|\nabla\hat{u}\|.$
Using the bound (6)
$b(\hat{u},u^{k}-u,\hat{u})$ $\geq-C_{s}\|\nabla\hat{u}\|^{2}\|\nabla(u^{k}-u)\|,$
$\hat{b}(\hat{u},u,\hat{u})$ $\geq-C_{s}\|\nabla\hat{u}\|^{2}\|\nabla u\|\geq-C_{s}M_{1}\|\nabla\hat{u}\|^{2},$
$b(\hat{u},T^{k}-T,\hat{T})$ $\geq-C_{s}\|\nabla u\|\nabla(T^{k}-T)\|\|\nabla\hat{T}\|,$
$\hat{b}(\hat{u},T,\hat{T})$ $\geq-C_{s}\|\nabla u\|\|\nabla T\|\|\nabla\hat{T}\|\geq-C_{s}M_{2}\|\nabla\hat{u}\|\|\nabla\hat{T}\|,$
Then we bound $A$ using the same bounds as above to get
$A((\hat{u},T),(\hat{u},T))$ $\geq(\nu-C_{s}\|\nabla(u^{k}-u)\|-C_{s}M_{1})\|\nabla\hat{u}\|^{2}$
$~{}~{}~{}-(C_{s}\|\nabla(T^{k}-T)\|+C_{s}M_{2}+C_{p}^{2}R_{i})\|\nabla u\|\|\nabla\hat{T}\|+\kappa\|\nabla\hat{T}\|^{2}$
$\geq(\nu-C_{s}\|\nabla(u^{k}-u)\|-C_{s}M_{1})\|\nabla\hat{u}\|^{2}$
$~{}~{}~{}-(C_{s}\|\nabla(T^{k}-T)\|+C_{s}M_{2}+C_{p}^{2}R_{i})\|\nabla\hat{u}\|^{2}$
$~{}~{}~{}-(\frac{C_{s}}{4}\|\nabla(T^{k}-T)\|+\frac{C_{s}}{4}M_{2}+\frac{C_{p}^{2}}{4}R_{i})\|\nabla\hat{T}\|^{2}+\kappa\|\nabla\hat{T}\|^{2}$
$\geq(\nu-(C_{s}\|\nabla(u^{k}-u)\|+C_{s}M_{1}+C_{s}\|\nabla(T^{k}-T)\|+C_{s}M_{2}+C_{p}^{2}R_{i}))\|\nabla\hat{u}\|^{2}$
$~{}~{}~{}+(\kappa-(\frac{C_{s}}{4}\|\nabla(T^{k}-T)\|+\frac{C_{s}}{4}M_{2}+\frac{C_{p}^{2}}{4}R_{i}))\|\nabla\hat{T}\|^{2}$
$\geq\sigma_{min}\|(\hat{u},\hat{T})\|_{Y}^{2},$
where
$\sigma_{min}=\min\{(\nu-(C_{s}\|\nabla(u^{k}-u)\|+C_{s}M_{1}+C_{s}\|\nabla(T^{k}-T)\|+C_{s}M_{2}+C_{p}^{2}R_{i})),\\ (\kappa-(\frac{C_{s}}{4}\|\nabla(T^{k}-T)\|+\frac{C_{s}}{4}M_{2}+\frac{C_{p}^{2}}{4}R_{i}))\}.$
Hence A is coercive. Thus by Lax-Milgram, the Newton iteration (14) is well-posed.
With this well-posedness result, the solution operator of (14) defined by $g_{N}:(V,D)\rightarrow(V,D)$, $g_{N}(u^{k},T^{k})=(u^{k+1},T^{k+1})$ is well-defined provided the data restrictions of Lemma 2.6 are satisfied. We note this is only a sufficient condition, and the Newton iteration for the non-isothermal flow system (14) is believed to be well-posed on a much less restrictive set of parameters.
Lemma 2.7.
Assume for any $k$ that
$\gamma_{k}:=\alpha_{1}+\alpha_{2}+\nu^{-1}C_{s}\|\nabla(u-u^{k})\|+\kappa^{-1}C_{s}\|\nabla(T-T^{k})\|<1,$
and
$C_{p}^{2}\nu^{-1}R_{i}<1.$
Then the Newton iteration (5) converges quadratically for any $u^{k}$ and $T^{k}$ satisfying
$C_{s}(1-\max\{\gamma_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}(\kappa^{-1}+\nu^{-1})(\|\nabla(u-u^{k})\|+\|\nabla(T-T^{k})\|)<1.$
Proof.
Let $e_{T}^{k+1}=T-T^{k+1}$ and $e^{k}=u-u^{k}$. We subtract (8) from (14), add and subtract $b(u,e^{k+1},v)$ in the first equation and $\hat{b}(u^{k+1},T,v)$ into the second, and set $v=e^{k+1}$ and $w=e_{T}^{k+1}$ to get
$\begin{cases}b(e^{k+1},e^{k},e^{k+1})&+b(e^{k+1},u,e^{k+1})+b(e^{k},e^{k+1},e^{k+1})+b(u,e^{k+1},e^{k+1})+\nu\|\nabla e^{k+1}\|^{2}\\ &=R_{i}((0~{}e_{T}^{k+1})^{T},e^{k+1})+b(e^{k},e^{k},e^{k+1}),\\ \hat{b}(e^{k+1},e_{T}^{k},e_{T}^{k+1})&+\hat{b}(e^{k+1},T,e_{T}^{k+1})+\hat{b}(e^{k},e_{T}^{k+1},e_{T}^{k+1})+\hat{b}(u,e_{T}^{k+1},e_{T}^{k+1})+\kappa\|\nabla e_{T}^{k+1}\|^{2}\\ &=\hat{b}(e^{k},e_{T}^{k},e_{T}^{k+1}).\end{cases}$
Next using skew symmetry, 4 terms in the system above vanish, leaving us with
$\begin{cases}\nu\|\nabla e^{k+1}\|^{2}&=R_{i}((0~{}e_{T}^{k+1})^{T},e^{k+1})+b(e^{k},e^{k},e^{k+1})-b(e^{k+1},e^{k},e^{k+1})-b(e^{k+1},u,e^{k+1}),\\ \kappa\|\nabla e_{T}^{k+1}\|^{2}&=\hat{b}(e^{k},e_{T}^{k},e_{T}^{k+1})-\hat{b}(e^{k+1},e_{T}^{k},e_{T}^{k+1})-\hat{b}(e^{k+1},T,e_{T}^{k+1}).\end{cases}$
We use the bound (6), Cauchy-Schwarz, Poincaré, and the stability bounds from Lemma 2.1 to upper bound the right hand side terms as
$\begin{cases}\nu\|\nabla e^{k+1}\|^{2}&\leq C_{p}^{2}R_{i}\|\nabla e_{T}^{k+1}\|\|\nabla e^{k+1}\|+C_{s}\|\nabla e^{k}\|^{2}\|\nabla e^{k+1}\|+C_{s}\|\nabla e^{k+1}\|^{2}\|\nabla e^{k}\|+\nu\alpha_{1}\|\nabla e^{k+1}\|^{2},\\ \kappa\|\nabla e_{T}^{k+1}\|^{2}&\leq C_{s}\|\nabla e^{k}\|\|\nabla e_{T}^{k}\|\|\nabla e_{T}^{k+1}\|+C_{s}\|\nabla e^{k+1}\|\|\nabla e_{T}^{k}\|\|\nabla e_{T}^{k+1}\|+\kappa\alpha_{2}\|\nabla e^{k+1}\|\|\nabla e_{T}^{k+1}\|.\end{cases}$
This reduces to
$\begin{cases}(1-\alpha_{1}-\nu^{-1}C_{s}\|\nabla e^{k}\|)\|\nabla e^{k+1}\|&\leq C_{p}^{2}\nu^{-1}R_{i}\|\nabla e_{T}^{k+1}\|+C_{s}\nu^{-1}\|\nabla e^{k}\|^{2},\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\|\nabla e_{T}^{k+1}\|&\leq\kappa^{-1}C_{s}\|\nabla e^{k}\|\|\nabla e_{T}^{k}\|+\kappa^{-1}C_{s}\|\nabla e^{k+1}\|\|\nabla e_{T}^{k}\|+\alpha_{2}\|\nabla e^{k+1}\|.\end{cases}$ (15)
Next we add the equations and reduce to get
$(1-\alpha_{1}-\alpha_{2}-\nu^{-1}C_{s}\|\nabla e^{k}\|-\kappa^{-1}C_{s}\|\nabla e_{T}^{k}\|)\|\nabla e^{k+1}\|$ $+(1-C_{p}^{2}\nu^{-1}R_{i})\|\nabla e_{T}^{k+1}\|$
$\leq C_{s}\nu^{-1}\|\nabla e^{k}\|^{2}+\kappa^{-1}C_{s}\|\nabla e^{k}\|\|\nabla e_{T}^{k}\|,$
$\leq C_{s}(\nu^{-1}+\kappa^{-1})(\|\nabla e^{k}\|+\|\nabla e_{T}^{k}\|)^{2}.$
The left hand side is lower bounded as follows,
$(1-\max\{\gamma_{k},C_{p}^{2}R_{i}\nu^{-1}\})(\|\nabla e^{k+1}\|+\|\nabla e_{T}^{k+1}\|)\leq(1-\alpha_{1}-\alpha_{2}-\nu^{-1}C_{s}\|\nabla e^{k}\|-\kappa^{-1}C_{s}\|\nabla e_{T}^{k}\|)\|\nabla e^{k+1}\|\\ +(1-C_{p}^{2}\nu^{-1}R_{i})\|\nabla e_{T}^{k+1}\|.$
This reduces to,
$\|\nabla e^{k+1}\|+\|\nabla e_{T}^{k+1}\|\leq C_{s}(1-\max\{\gamma_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}(\kappa^{-1}+\nu^{-1})(\|\nabla e^{k}\|+\|\nabla e_{T}^{k}\|)^{2},$
which completes the proof.
3 Nonlinear Preconditioner Analysis
In this section, we consider (AA-)Picard-Newton for the non-isothermal flow system, then analyze its stability and convergence properties. We show that preconditioning stabilizes Newton and reduces the sufficient conditions on small data assumptions and closeness assumption of the initial guess. We restrict our analysis to the case of no AA, however we note that with extra technicalities, similar proofs could be adapted.
Using the solution operators for Picard and Newton defined in the previous section, the Picard based nonlinear preconditioning of Newton, Picard-Newton, is defined as $g_{PN}:=g_{N}\circ g_{P}$ so that
$g_{PN}(u^{k},T^{k})=g_{N}(g_{P}(u^{k},p^{k},T^{k}))=(u^{k+1},T^{k+1}).$
Thus, Picard-Newton consists of two steps. The first step is the Picard step given by $g_{P}(u^{k},T^{k})=(\tilde{u}^{k+1},\tilde{T}^{k+1})$ which is then followed by the Newton step given by $g_{N}(\tilde{u}^{k+1},\tilde{T}^{k+1})=(u^{k+1},T^{k+1})$.
Algorithm 3.1.
Picard-Newton
Given $(u^{0},p^{0},T^{0})\in(X,Q,D)$.
for $k=0,1,2,...$ do
     Step 1: Solve for $(\tilde{u}^{k+1},\tilde{T}^{k+1})\in V\times D$ satisfying
$\begin{cases}b(u^{k},\tilde{u}^{k+1},v)+\nu(\nabla\tilde{u}^{k+1},\nabla v)&=(f,v)+R_{i}((0~{}\tilde{T}^{k+1})^{T},v),\\ \hat{b}(u^{k},\tilde{T}^{k+1},w)+\kappa(\nabla\tilde{T}^{k+1},\nabla w)&=(g,w).\end{cases}$ (16)
     Step 2: Solve for $(u^{k+1},T^{k+1})\in V\times D$ satisfying
$\begin{cases}b(u^{k+1},\tilde{u}^{k+1},v)+b(\tilde{u}^{k+1},u^{k+1},v)+\nu(\nabla u^{k+1},\nabla v)&=(f,v)+R_{i}((0~{}T^{k+1})^{T},v)\\ &~{}~{}~{}+b(\tilde{u}^{k+1},\tilde{u}^{k+1},v),\\ \hat{b}(u^{k+1},\tilde{T}^{k+1},w)+\hat{b}(\tilde{u}^{k+1},T^{k+1},w)+\kappa(\nabla T^{k+1},\nabla w)&=(g,w)+\hat{b}(\tilde{u}^{k+1},\tilde{T}^{k+1},w).\end{cases}$ (17)
end for
Picard is known to be unconditionally stable, while Newton is not (and is consequently divergent for many $Ra$). Therefore it is intuitive that the stability of Picard will have a positive effect on the stability of Picard-Newton. Hence, we begin by analyzing the stability of Picard-Newton.
Theorem 3.2.
If $\alpha_{1}+\alpha_{2},~{}\nu^{-1}R_{i}C_{p}^{2}<1$, Picard-Newton (16)-(17) is stable with
$(1-\alpha_{1}-\alpha_{2})\|\nabla u^{k+1}\|+(1-\nu^{-1}R_{i}C_{p})\|\nabla T^{k+1}\|\leq M_{2}+2M_{1}.$
Remark.
The small data conditions for stability of Picard-Newton are the same as the uniqueness condition for the non-isothermal flow system.
Proof.
We begin by considering the stability bound resulting from the Picard step. This part of the proof follows analogously to Lemma 2.3 resulting in
$\begin{cases}\|\nabla\tilde{u}^{k+1}\|&\leq M_{1},\\ \|\nabla\tilde{T}^{k+1}\|&\leq M_{2}.\end{cases}$
Next we consider (17) and choose $v=u^{k+1}$ and $w=T^{k+1}$. This gives
$\begin{cases}\nu\|\nabla u^{k+1}\|^{2}&=(f,u^{k+1})+R_{i}((0~{}T^{k+1})^{T},u^{k+1})+b(\tilde{u}^{k+1},\tilde{u}^{k+1},u^{k+1})-b(u^{k+1},\tilde{u}^{k+1},u^{k+1}),\\ \kappa\|\nabla T^{k+1}\|^{2}&=(g,T^{k+1})+\hat{b}(\tilde{u}^{k+1},\tilde{T}^{k+1},T^{k+1})-\hat{b}(u^{k+1},\tilde{T}^{k+1},T^{k+1}).\end{cases}$
We next use Cauchy-Schwarz, upper bound the trilinear term, use Poincaré, and reduce to get
$\begin{cases}\nu\|\nabla u^{k+1}\|&\leq\|f\|_{V^{\ast}}+R_{i}C_{p}^{2}\|\nabla T^{k+1}\|+C_{s}\|\nabla\tilde{u}^{k+1}\|^{2}+C_{s}\|\nabla u^{k+1}\|\|\nabla\tilde{u}^{k+1}\|,\\ \kappa\|\nabla T^{k+1}\|&\leq\|g\|_{D^{\ast}}+C_{s}\|\nabla\tilde{u}^{k+1}\|\|\nabla\tilde{T}^{k+1}\|+C_{s}\|\nabla u^{k+1}\|\|\nabla\tilde{T}^{k+1}\|.\end{cases}$
Using the stability bound for the Picard step and combining like terms
$\begin{cases}(1-\alpha_{1})\|\nabla u^{k+1}\|&\leq\nu^{-1}\|f\|_{V^{\ast}}+\nu^{-1}R_{i}C_{p}^{2}\|\nabla T^{k+1}\|+\alpha_{1}M_{1},\\ \|\nabla T^{k+1}\|&\leq\kappa^{-1}\|g\|_{D^{\ast}}+\alpha_{2}M_{1}+\alpha_{2}\|\nabla u^{k+1}\|,\end{cases}$
where $\alpha_{1}=\nu^{-1}C_{s}M_{1}$ and $\alpha_{2}=\kappa^{-1}C_{s}M_{2}$.
We add the equations and combine like terms to obtain
$(1-\alpha_{1}-\alpha_{2})\|\nabla u^{k+1}\|+(1-\nu^{-1}R_{i}C_{p})\|\nabla T^{k+1}\|\leq\nu^{-1}\|f\|_{V^{\ast}}+\kappa^{-1}\|g\|_{D^{\ast}}+(\alpha_{1}+\alpha_{2})M_{1}.$
Using that $M_{2}=\kappa^{-1}\|g\|_{D^{\ast}}$ and $M_{1}=\nu^{-1}\|f\|_{V^{\ast}}+R_{i}C_{p}^{2}\nu^{-1}M_{2}$ we can upper bound the left hand side to get the bound
$(1-\alpha_{1}-\alpha_{2})\|\nabla u^{k+1}\|+(1-\nu^{-1}R_{i}C_{p})\|\nabla T^{k+1}\|$ $\leq M_{2}+(\alpha_{1}+\alpha_{2}+1)M_{1},$ $\leq M_{2}+2M_{1}.$
One way to understand the effect of Picard’s stability on Newton is to consider the case when $u^{k}$ does not satisfy the sufficient conditions for convergence of Newton. In this case, one would expect that the Newton step could produce solutions which are divergent. However this is not the case because of Picard’s stability bound.
Picard-Newton is a two step method and so for clarity we split the error analysis into two theorems: a theorem for the error arising in the Picard step and a theorem for the iterations error after the Newton step. The analysis for the Picard step will follow very closely to the analysis for the usual Picard iteration and thereby includes the constants $M_{1}$ and $M_{2}$ that naturally arise in the analysis.
Theorem 3.3.
Assume $C_{p}^{2}Ri\nu^{-1}<1$. Then the Picard step (16) of Picard-Newton satisfies
$\|\nabla(u-\tilde{u}^{k+1})\|+(1-R_{i}C_{p}^{2}\nu^{-1})\|\nabla(T-\tilde{T}^{k+1})\|\leq(\alpha_{1}+\alpha_{2})\|\nabla(u-u^{k})\|.$
In particular we have,
$\|\nabla(T-\tilde{T}^{k+1})\|\leq\alpha_{2}\|\nabla(u-u^{k})\|,$ (18)
and
$\|\nabla(u-\tilde{u}^{k+1})\|\leq(\alpha_{1}+\alpha_{2})\|\nabla(u-u^{k})\|.$ (19)
Proof.
This proof follows analogously to Lemma 2.5.
Note that the bounds above on $(u-\tilde{u}^{k+1})$ and $(T-\tilde{T}^{k+1})$ both depend upon $(u-u^{k})$. This means the accuracy of $\tilde{T}^{k+1}$ depends on the accuracy of $u^{k}$ for the Picard step.
Theorem 3.4.
Assume $R_{i}C_{p}^{2}\nu^{-1}<1$ and
$\beta_{k}:=\alpha_{1}+\alpha_{2}+C_{s}(\alpha_{1}+\alpha_{2})[\kappa^{-1}+\nu^{-1}]\|\nabla(u-u^{k})\|<1$
for any $k$. Then Picard-Newton converge quadratically for any $u^{k}$ satisfying
$(1-\max\{\beta_{k},R_{i}C_{p}^{2}\nu^{-1}\})^{-1}C_{s}(\alpha_{1}+\alpha_{2})^{2}(\nu^{-1}+\kappa^{-1})\|\nabla(u-u^{k})\|<1.$
Proof.
We subtract (17) from (8), then add and subtract terms and we the skew symmetry of the trilinear terms to write
$\begin{cases}b(e^{k+1},\tilde{e}^{k+1},e^{k+1})+b(e^{k+1},u,e^{k+1})+\nu\|\nabla e^{k+1}\|^{2}&=(R_{i}(0~{}e_{T}^{k+1})^{T},e^{k+1})+b(\tilde{e}^{k+1},\tilde{e}^{k+1},e^{k+1}),\\ \hat{b}(e^{k+1},\tilde{e}_{T}^{k+1},e_{T}^{k+1})+\hat{b}(e^{k+1},T,e_{T}^{k+1})+\kappa\|\nabla e_{T}^{k+1}\|^{2}&=\hat{b}(\tilde{e}^{k+1},\tilde{e}_{T}^{k+1},e_{T}^{k+1}),\end{cases}$
where $e_{T}^{k+1}=T^{k+1}-T$ and $e^{k+1}=u^{k+1}-u$.
We now use the bounds from Lemma 2.1 to get the upper bounds
$\begin{cases}\nu\|\nabla e^{k+1}\|^{2}&\leq C_{p}^{2}R_{i}\|\nabla e_{T}^{k+1}\|\|\nabla e^{k+1}\|+C_{s}\|\nabla\tilde{e}^{k+1}\|^{2}\|\nabla e^{k+1}\|+C_{s}\|\nabla e^{k+1}\|^{2}\|\nabla\tilde{e}^{k+1}\|+\alpha_{1}\nu\|\nabla e^{k+1}\|^{2},\\ \kappa\|\nabla e_{T}^{k+1}\|^{2}&\leq C_{s}\|\nabla\tilde{e}^{k+1}\|\|\nabla\tilde{e}_{T}^{k+1}\|\|\nabla e_{T}^{k+1}\|+C_{s}\|\nabla e^{k+1}\|\|\nabla\tilde{e}_{T}^{k+1}\|\|\nabla e_{T}^{k+1}\|+\kappa\alpha_{2}\|\nabla e^{k+1}\|\|\nabla e_{T}^{k+1}\|,\end{cases}$
which reduces to
$\begin{cases}(1-\alpha_{1})\|\nabla e^{k+1}\|&\leq C_{p}^{2}\nu^{-1}R_{i}\|\nabla e_{T}^{k+1}\|+C_{s}\nu^{-1}\|\nabla\tilde{e}^{k+1}\|^{2}+C_{s}\nu^{-1}\|\nabla e^{k+1}\|\|\nabla\tilde{e}^{k+1}\|,\\ \|\nabla e_{T}^{k+1}\|&\leq C_{s}\kappa^{-1}\|\nabla\tilde{e}^{k+1}\|\|\nabla\tilde{e}_{T}^{k+1}\|+C_{s}\kappa^{-1}\|\nabla e^{k+1}\|\|\nabla\tilde{e}_{T}^{k+1}\|+\alpha_{2}\|\nabla e^{k+1}\|.\end{cases}$
Next we continue bounding using (19) and (18) to obtain
$\begin{cases}(1-\alpha_{1})\|\nabla e^{k+1}\|&\leq C_{p}^{2}\nu^{-1}R_{i}\|\nabla e_{T}^{k+1}\|+C_{s}\nu^{-1}(\alpha_{1}+\alpha_{2})^{2}\|\nabla e^{k}\|^{2}+C_{s}\nu^{-1}(\alpha_{1}+\alpha_{2})\|\nabla e^{k+1}\|\|\nabla e^{k}\|,\\ \|\nabla e_{T}^{k+1}\|&\leq C_{s}\kappa^{-1}\alpha_{2}(\alpha_{1}+\alpha_{2})\|\nabla e^{k}\|^{2}+C_{s}\kappa^{-1}\alpha_{2}\|\nabla e^{k+1}\|\|\nabla e^{k}\|+\alpha_{2}\|\nabla e^{k+1}\|.\end{cases}$
Adding the equations and reducing yields
$(1-\alpha_{1}-\alpha_{2}-C_{s}[\kappa^{-1}\alpha_{2}$ $+\nu^{-1}(\alpha_{1}+\alpha_{2})]\|\nabla e^{k}\|)\|\nabla e^{k+1}\|+(1-C_{p}^{2}\nu^{-1}R_{i})\|\nabla e_{T}^{k+1}\|$
$\leq C_{s}(\alpha_{1}+\alpha_{2})(\nu^{-1}(\alpha_{1}+\alpha_{2})+\kappa^{-1}\alpha_{2})\|\nabla e^{k}\|^{2}$
$\leq C_{s}(\alpha_{1}+\alpha_{2})^{2}(\nu^{-1}+\kappa^{-1})\|\nabla e^{k}\|^{2}.$
We now lower bound the left hand side as
$(1-\max\{\beta_{k},C_{p}^{2}R_{i}\nu^{-1}\})(\|\nabla e^{k+1}\|+\|\nabla e_{T}^{k+1}\|)\leq(1-\alpha_{1}-\alpha_{2}-C_{s}[\kappa^{-1}\alpha_{2}+\nu^{-1}(\alpha_{1}+\alpha_{2})]\|\nabla e^{k}\|)\|\nabla e^{k+1}\|\\ +(1-C_{p}^{2}R_{i}\nu^{-1})\|\nabla e_{T}^{k+1}\|,$
and upper bound the right hand side to obtain
$\|\nabla e^{k+1}\|+\|\nabla e_{T}^{k+1}\|\\ \leq(1-\max\{\beta_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}C_{s}(\alpha_{1}+\alpha_{2})^{2}(\nu^{-1}+\kappa^{-1})\|\nabla e^{k}\|^{2}.$
We provide table 1 below for a more clear comparison of the sufficient conditions for Newton and Picard-Newton. First notice that the assumptions for Picard-Newton include similar terms, and in fact they both require $C_{p}^{2}R_{i}\nu^{-1}<1$. However, it is clear that Picard-Newtons assumptions are less restrictive as they do not include any assumptions on $T$, and the velocity term in $\beta_{k}$ in Picard-Newton is scaled by $(\alpha_{1}+\alpha_{2})$ compared to Newton (note that this same condition ensures $(\alpha_{1}+\alpha_{2})<1$). These weakened restriction for Picard-Newton are due to the Picard step allowing $u^{k}$ to control the error of $T$.
Similarly, the sufficient condition on the initial guess for convergence of Picard-Newton is scaled by $(\alpha_{1}+\alpha_{2})^{2}$ compared to Newton. Because another sufficient condition for Picard-Newton ensures that $\alpha_{1}+\alpha_{2}<1$, this allows more $u^{k}$ to satisfy the equation
$C_{s}(1-\max\{\beta_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}(\alpha_{1}+\alpha_{2})^{2}(\nu^{-1}+\kappa^{-1})\|\nabla(u-u^{k})\|<1.$
However there is another important connection to be made between Picard and the sufficient conditions. If we consider just Picard, it is required for its convergence that $\alpha_{1}+\alpha_{2}<1$. Then for Picard-Newton, $\alpha_{1}$ and $\alpha_{2}$ come from the error of the Picard step and appear in the sufficient conditions and assumptions of the method. This means that Picard-Newtons conditions for convergence are potentially being aided by Picard’s convergence properties. Furthermore, recall that the assumption for uniqueness of solutions to non-isothermal flow system includes $\alpha_{1}+\alpha_{2}<1$. This further shows the assumption is reasonable and therefore Picard-Newton’s sufficient conditions for convergence of Picard-Newton are improved when compared to Newton.
Sufficient conditions
Newton
$\gamma_{k}=\alpha_{1}+\alpha_{2}+\nu^{-1}C_{s}\|\nabla(u-u^{k})\|+\kappa^{-1}C_{s}\|\nabla(T-T^{k})\|<1$
$C_{p}^{2}\nu^{-1}R_{i}<1$
$(1-\max\{\gamma_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}(\kappa^{-1}+\nu^{-1})(\|\nabla(u-u^{k})\|+\|\nabla(T-T^{k})\|)<1$
Picard-Newton
$\beta_{k}=\alpha_{1}+\alpha_{2}+C_{s}(\alpha_{1}+\alpha_{2})[\kappa^{-1}+\nu^{-1}]\|\nabla(u-u^{k})\|<1$
$C_{p}^{2}R_{i}\nu^{-1}<1$
$(1-\max\{\beta_{k},C_{p}^{2}R_{i}\nu^{-1}\})^{-1}(\alpha_{1}+\alpha_{2})^{2}(\nu^{-1}+\kappa^{-1})\|\nabla(u-u^{k})\|<1$
Table 1: Shown above is a comparison of sufficient conditions for Newton and Picard-Newton convergence.
4 Numerical Results
We give now numerical results for two test problems for (AA-)Picard-Newton applied to the non-isothermal flow system. We consider the non-isothermal flow system with $g=0$, $f=0$, $\nu=\kappa=10^{-1}$, and $Ri$ is varied which varies $Ra$ for the problems. The domain $\Omega$ and boundary conditions are different for each test and will therefore be provided in the following subsections.
Let $\tau_{h}$ be a barycentric refined triangular mesh for $\Omega$. We use $X_{h}=\mathbb{P}_{2}(\tau_{h})\cap X$, $Q_{h}=\mathbb{P}_{1}^{disc}(\tau_{h})\cap Q$, and $D_{h}=\mathbb{P}_{2}(\tau_{h})\cap D$. We note that $(X_{h},Q_{h})$ are LBB stable in this setting, and also provide divergence free velocity solutions [1], so all the analysis of the previous sections applies. We also choose the initial guess to be $(u_{0},p_{0},T_{0})=(0,0,0)$. Using the B-norm defined below, the convergence criteria will be given as
$\|(u^{k},T^{k})-g_{PN}(u^{k},T^{k})\|_{B}:=\sqrt{\nu\|\nabla(u^{k}-u^{k-1})\|^{2}+\kappa\|\nabla(T^{k}-T^{k-1})\|^{2}}<10^{-8},$
and we use 200 iterations as the maximum number of allowed iterations.
4.1 Differentially Heated Cavity
The first numerical test is the classical differentially heated cavity problem [3]. The domain is the unit square $\Omega=(0,1)^{2}$, with boundary conditions
$\begin{cases}u&=0~{}\text{ on }\partial\Omega,\\ T(0,y)&=0,\\ T(1,y)&=1,\\ \nabla T\cdot n&=0,\text{ on}~{}x=0,~{}x=1.\end{cases}$
We use a mesh with max element diameter $h=1/64$ and note that all results are comparable for other choices of h that we tested. Solutions to this system with $Ra=10000$ are shown in figure 1. We begin the study of the convergence of (AA-)Picard-Newton for varying $Ra$ by considering Picard-Newton and then AA-Picard-Newton.
4.1.1 Picard-Newton
Convergence results of Picard-Newton are shown in figure 2 for $Ra=50000$, $100000$, $150000$, $200000$, and $250000$. We observe that Picard-Newton converges for each $Ra$, and quadratic convergence is observed in the asymptotic range. As Ra increases, the number of iterations required for convergence increases as well. We observe the method remains stable even when it is (seemingly) not making progress in its residual.
Figure 1: Shown above is the computed solution of the differentially heated cavity problem for velocity streamlines (left) and temperature contours (right) for $Ra=10000$
For comparison, we also perform tests using Picard alone and Newton alone. In figure 3 (left) we see that Picard converges within 200 iterations for $Ra$ up to $10000$. For higher $Ra$, Picard does not converge, but it does remain stable. Newton is able to converge for slightly higher $Ra$ than Picard, as shown in figure 3 (left): it converges for up to $Ra=15000$, but above this $Ra$ we observe divergence. Moreover, as is typical for Newton, when it diverges, it blows up. We observe then that that Picard-Newton is able to converge for $Ra$ a full order of magnitude higher than Picard and Newton.
Figure 2: Shown above are convergence plots for Picard-Newton at varying $Ra$.
Figure 3: Shown above are convergence plots for Picard (left) and Newton(right) at varying $Ra$.
4.1.2 AA-Picard-Newton for Differentially Heated Cavity
We now apply AA-Picard-Newton to the differentially heated cavity problem. Recall that for this problem Picard-Newton converges for $Ra$ as high as $250000$. Convergence for AA-Picard-Newton are shown in figure 4 for varying $Ra$ and $m=1$ (left) and $m=3$ (right). AA-Picard-Newton shows significant improvement in convergence for higher $Ra$ both in iteration count and solvability. Convergence is even further improved with increased depth: for $m=1$ we observe convergence up to $Ra=500000$ and with $m=3$ convergence is obtained up to $Ra=750000$ (and convergence is accelerated by $m=3$ over $m=1$ for $Ra$ where $m=1$ converges).
Method Max $Ra$ for convergence
AA-Picard-Newton: m=3 $750000$
AA-Picard-Newton: m=1 $500000$
Picard-Newton $250000$
Picard $10000$
Newton $15000$
Table 2: Shown above is the maximum $Ra$ (to the nearest 1000) for which each method converges.
Figure 4: Convergence plots for AA-Picard-Newton with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
4.2 Heated Cavity on a Complex Domain
The second numerical test is on a more complex domain shown in figure 5, with boundary conditions
$\begin{cases}u&=0~{}~{}\text{on }\partial\Omega,\\ T(x,1)&=1,\\ T(x,0)&=\frac{2x}{7},\\ \nabla T\cdot n&=0~{}~{}0<y<1.\end{cases}$
Solutions for $u$ and $T$ shown on the domain in figure 5 for $Ra=1000$. We use a Delaunay mesh with max element diameter $h=7/64$, and again note that all results are comparable for other mesh sizes that we tested. We again begin the study of the convergence of (AA-)Picard-Newton for this heated cavity problem by considering Picard-Newton and then AA-Picard-Newton.
Figure 5: Shown above is the computed solution for $u$ (left) and $T$ (right) for $Ra=1000$ in the second numerical test.
4.2.1 Picard-Newton
The convergence results are shown in figure 6, for $Ra$ $=$ $5000$, $10000$, $15000$, $20000$, $25000$, and $50000$. As in the previous test problem, we observe quadratic convergence for each $Ra$ which Picard-Newton converges, for (up to $Ra=10000$). For $Ra>10000$, we do not observe convergence within 200 iterations but note the computed solutions remain stable.
Again for comparison, we solve the test problem using both Picard alone and Newton alone, and display their convergence results in figure 7. We see that Picard converges linearly for $Ra$ up to $1000$. For higher $Ra$, Picard does not converge but remains stable. Newton, in figure 7 (left), converges for $Ra$ up to $2500$ which is slightly higher than Picard. For $Ra=5000$, convergence is not achieved but Newton curiously remains stable. For $Ra>5000$ we see divergence, and solutions blow up. In comparison, using Picard-Newton we get convergence for $Ra=20000$ with solutions that remain stable for all $Ra$ that we tested.
Figure 6: Shown above are convergence plots for Picard-Newton at varying $Ra$.
Figure 7: Shown above are convergence plots for Picard (left) and Newton (right) at varying $Ra$.
4.2.2 AA-Picard-Newton on a Complex Domain
We next apply AA-Picard-Newton to the heated cavity problem on a complex domain from section 4.2. Recall that for this problem (AA-)Picard-Newton converges in 200 iterations for $Ra$ up to $10000$. AA-Picard-Newton achieves convergence with depth $m=1$ for $Ra$ up to $125000$ as shown in figure 8. This is further improved with depth $m=3$ where AA-Picard-Newton converges for $Ra$ up to $75000$. As before, for higher $Ra$ we see stability for the method and it may converge if allowed to continue past 200 iterations.
Method Max $Ra$ for convergence
AA-Picard-Newton: m=3 $75000$
AA-Picard-Newton: m=1 $125000$
Picard-Newton $20000$
Picard $1000$
Newton $5000$
Table 3: Shown above is the maximum $Ra$ (to the nearest 100) for which each method converges.
Figure 8: Shown above are convergence plots for AA-Picard-Newton convergence with $m=1$ (left) and $m=3$ (right) at varying $Ra$.
5 Conclusion
(AA-)Picard-Newton for the non-isothermal flow system is an improvement upon the Newton iteration that provides less restrictive (sufficient) conditions for quadratic convergence, and dramatically better convergence results on the benchmark problems. The sufficient conditions of Picard-Newton incorporate the accuracy bounds of the Picard step, which leads to a scaling of the sufficient conditions of Newton by a constant less than $1$. Furthermore, the sufficient conditions of Picard-Newton depend solely upon the fluid velocity $u^{k}$ due to the preconditioning by Picard causing accuracy of $T^{k}$ to depend on $u^{k}$. This is in stark contrast to the usual Newton whose sufficient conditions depends upon both $u^{k}$ and $T^{k}$. Furthermore, Picard-Newton is stable for any initial guess, provided the data satisfy a smallness condition that also implies uniqueness of the PDE system. These analytical results are extendable to AA-Picard-Newton with extra technical difficulties.
This improved theory is reflected in the numerical tests where for a simple domain we see convergence for high $Ra$ and errors that remain stable. In comparison, Newton diverges for relatively low $Ra$ and Picard converges slower for the same $Ra$. This is further shown in the non-isothermal flow system on a complex domain which is comparatively harder. (AA-)Picard-Newton converges for high $Ra$ and is proved stable, while Picard and Newton either converges slower or fail to converge for these same $Ra$.
6 Acknowledgements
The author was partially supported by NSF grant DMS 2011490.
The author thanks Professor Leo Rebholz for helpful discussions regarding this work.